OpenStudy (anonymous):

Two different radioactive isotopes decay to 10% of their respective original amounts. Isotope A does this is 33 days, while isotope B does this in 43 days. What is the approximate difference in the half-lives of the isotopes? 3 days 10 days 13 days 33 days

OpenStudy (anonymous):

@Nnesha

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i guess you need the half life of each right?

OpenStudy (anonymous):

yeah

OpenStudy (misty1212):

i am trying to think of a quick way to do it, without going through the whole \[A_0e^{kt}\] business give me a sec, otherwise we have to grind it out

OpenStudy (misty1212):

yeah we can do it a quicker way solve \[\left(\frac{1}{10}\right)^{\frac{t}{33}}=\frac{1}{2}\]for \(t\) to find the half life of the first one for the second solve \[\left(\frac{1}{10}\right)^{\frac{t}{43}}=\frac{1}{2}\]

OpenStudy (anonymous):

so t would be the half life?

OpenStudy (misty1212):

when you solve that, yes you want to do it the hard way or the easy way?

OpenStudy (anonymous):

Nope, just solved it in the calculator... it ended up being a difference of three days :)

OpenStudy (anonymous):

Thanks so much

OpenStudy (misty1212):

ok good

OpenStudy (misty1212):

yeah first close to 10, second to 13 \[\color\magenta\heartsuit\]

OpenStudy (anonymous):

yep, much faster than the normal formula